# MNSC-0001-2024.00

### 15 Mar 2026 v1

@@ 4.1 Model Setup
Balance sheet information.
We consider a financial system comprised of $m$ banks, indexed by $1,\dots,m$, that operates over two periods ($t=1,2$). Each bank $i$’s balance sheet consists of the following four components:
@@
  • $z_{i}\geq 0$ represents the “volume” of its illiquid assets, with a nominal unit price of PP; [^3] We adopt the standard assumption in the literature that all banks hold the same type of illiquid assets, which serves as an approximation of multiple correlated assets (WeberWeske2017, BraouezecWagalath2019).
@@
We assume that these components remain constant throughout the two periods and that liquid assets (e.g., cash, central bank reserves, and short-term government securities) can be sold and transferred at their nominal value without any cost. To exclude trivial cases where banks are not required to liquidate their illiquid assets, it is assumed that $e_{i}<L_{i}$ for all $i\in[m]$. The associated balance sheet is described in Table 4.1.
[…]
Initial balance sheet of bank $i$ before liquidation Assets Liabilities and Net Worth Liquid assets $e_{i}$ Liabilities $L_{i}$ Illiquid assets $z_{i}$ with price $P$ Net worth $w_{i}$
Illiquid asset price modeling.
@@
Illiquid assets (e.g., loans, mortgages, and real estate holdings) are, however, difficult to sell without a significant loss of value. In the related literature (CifuentesEtAl2005, ChenEtAl2016), this loss is typically modeled by the so-called inverse demand function $Q:[0,\sum_{j=1}^{m}z_{j}]\rightarrow\mathbb{R}_{+}$, a decreasing function that maps the aggregate supply of illiquid assets across all banks to their price. Throughout this paper, we impose several regularity conditions on the function $Q$, described below: {assumption}The inverse demand function $Q$ is twice continuously differentiable and satisfies the following five conditions: (i) $Q(0)=P$; (ii) $Q^{\prime}(x)<0$; (iii) $(xQ(x))^{\prime}>0$; (iv) $Q^{\prime\prime}(x)\geq 0$; and (v) $(xQ^{\prime}(x))^{\prime}<0$. The first condition states that illiquid assets are valued at their nominal price in the absence of sales. Conditions (ii) and (iii) reflect the observation that as the supply of illiquid assets increases, their price decreases while the total value of these assets in the market grows. The fourth condition imposes a diminishing marginal effect on the price of illiquid assets with respect to their size. Condition (v) ensures that the total value of illiquid assets in the market is concave in their size. Note that the above conditions (or their variants) are commonly used in the literature (AminiEtAl2016c, BichuchFeinstein2019). Examples of inverse demand functions satisfying Assumption 4.1 include Q(x)=PaxQ(x)=P-ax with $0<a<P/(2\sum_{k=1}^{m}z_{k})$, $Q(x)=P\exp(-ax)$ with $0<a<1/\sum_{k=1}^{m}z_{k}$, and $Q(x)=P(\beta x+1)^{-a}$ with $0<a<1/(\beta\sum_{k=1}^{m}z_{k})$ and $\beta>0$.
Early liquidation and permanent price impact.
@@
We assume that all banks clear their debts at the end of the time horizon (i.e., $t=2$), consistent with prior studies (e.g., CifuentesEtAl2005, ChenEtAl2016, AminiEtAl2016c, GhamamiEtAl2022). As alluded to earlier, a salient feature of our modeling framework lies in the option for strategic early liquidation at $t=1$ to mitigate liquidation losses and to sell a smaller volume of illiquid assets, which is in stark contrast to most existing models that allow liquidation only at clearing. Suppose that each bank $i$ chooses to sell off $x_{i}$ units of illiquid assets at $t=1$. Then, $Q(\sum_{k=1}^{m}x_{k})$ becomes the market price of illiquid assets at $t=1$ and is strictly smaller than the nominal price $P$ by Assumption 4.1.
@@
To account for permanent price impact, we consider a situation where the price impact generated at $t=1$ persists into $t=2$. Specifically, the asset price at clearing is governed by another inverse demand function of the asset supply at $t=2$, defined as $u\mapsto Q(\gamma\sum_{k=1}^{m}x_{k}+u)$, where the parameter $\gamma\in[0,1]$ represents the proportion of the early liquidation volume that continues to depress prices at $t=2$. That is, a larger $\gamma$ implies slower price recovery. The limiting case $\gamma=0$ corresponds to fully transient impact, where the initial price at $t=2$ recovers to the nominal price $P$, as in the example in Section 3. Conversely, $\gamma=1$ signifies fully permanent price impact, where the opening price at $t=2$ remains at the post-liquidation price at $t=1$; this latter case coincides with the model in GhamamiEtAl2022. While existing research commonly focuses on either of these two extreme cases, our approach generalizes the framework of GhamamiEtAl2022 by allowing for “partially” permanent price impact via the parameter $\gamma$.
Clearing process at $t=2$.
@@
Following the early liquidation at $t=1$, each bank $i\in[m]$ retains $z_{i}-x_{i}$ units of illiquid assets, while its liquid asset value becomes $e_{i}+x_{i}Q(\sum_{k=1}^{m}x_{k})$. At the clearing stage, given a price $p$, each bank $i$ attempts to cover its liabilities by selling $p^{-1}(L_{i}-e_{i}-x_{i}Q(\sum_{k=1}^{m}x_{k}))^{+}$ units of illiquid assets. If the bank cannot meet the obligations, it defaults and liquidates all remaining $z_{i}-x_{i}$ units of assets. Thus, the quantity sold by bank $i$ is $\{p^{-1}(L_{i}-e_{i}-x_{i}Q(\sum_{k=1}^{m}x_{k}))^{+}\}\wedge(z_{i}-x_{i})$ for all $i\in[m]$. Consequently, the market price $p^{\gamma}_{2}$ of illiquid assets at $t=2$ is determined by the following fixed-point equation
@@
Note that a straightforward extension of Lemma 3 in AminiEtAl2016c ensures the uniqueness of $p^{\gamma}_{2}$. We denote each bank $i$’s liquidation size at $t=2$ by
@@
Remark 4.1
Permanent price impact could also be modeled using more sophisticated dynamics, such as, the Volume Weighted Average Price (VWAP) approach investigated in  BanerjeeFeinstein2021. Unlike static snapshots, this method captures the entire liquidation history by defining the average asset price through an integral of the inverse demand function. However, the technical analysis based on the VWAP framework introduces significant mathematical complexity due to the integration terms; therefore, we retain our current setup to preserve analytical tractability. We anticipate that the qualitative insights derived from our findings would persist under the VWAP framework.
4.2 Main Problem
@@
Robustification against other banks’ actions.
As shown in (6), each bank’s liquid asset value at clearing depends on the volume of illiquid assets sold by other banks at $t=1$, which is often unknown. Hence, the solution to the following problem yields bank $i$’s optimal early liquidation decision under the worst-case strategies of other banks:
@@
Relaxation of liquidation size at clearing.
Invoking Lemma 4.2, we circumvent the aforementioned intractability of (7) by relaxing the constraint on each bank $j$’s liquidation size at clearing. Specifically, we allow it to take any value in the interval $0,\bar{x}_{j}-x_{j}]$ rather than restricting it to $\psi_{j}({\bf x})$ for each $j\in[m]$. Under this relaxation, the maximin problem in (7) can be converted into the following form:
@@
For any feasible solution $({\bf x},{\bf y})$ of (8), Assumption 4.1 indicates that
@@
Validity of the formulation.
Before presenting the analysis of our problem (9), let us discuss its practical validity to address potential concerns about the implication of optimizing $\Phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$. Specifically, the following result provides insights into how the relaxation affects the liquidation strategy, or equivalently, the relationship between the solutions to (7) and (9)
Proposition 4.3 (Validity of (9))
@@
By the definitions of $\bar{x}_{i}$ and $\psi^{\gamma}_{j}({\bf x})$, it is easy to check that $\phi^{\gamma}_{i}(x_{i},\bar{{\bf x}}_{-i})\leq L_{i}$ for any $x_{i}\in[0,\bar{x}_{i}]$; see the proof of Proposition 4.3 for more details. [^5] From a practical perspective, this inequality holds since each bank’s liquid asset value at clearing is characterized as either exactly meeting the payment obligation or failing to do so. Thus, the first statement of Proposition 4.3 shows that if the optimal value of (9) is greater than or equal to $L_{i}$, then bank $i$’s robust optimal strategy, $x^{\gamma}_{i}$, is also an optimal solution to the outer problem of (7). Furthermore, the second statement of the theorem suggests that, even if the optimal value of (9) is smaller than $L_{i}$, bank $i$’s robust optimal strategy still maximizes a lower bound for the worst-case value of its liquid assets at clearing. Accordingly, both statements indicate that the robust optimal strategy obtained by solving (9) can serve as a conservative approximation and a reliable surrogate for the true optimal decision that solves (7). This answers our second research question (Q2), introduced in Section 3, for the case of permanent price impact.
@@
In Figure 4.2, we numerically validate the results in Proposition 4.3 by comparing the inner optimal values $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$ and $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\Phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$ across different values of $x_{i}$ in a two-bank system with $m=i=2$ for varying levels of permanent price impact: $\gamma=$ 0, $0.5$, and $1$. The former optimal value is plotted using blue curves—solid for $\gamma=0$, dashed for $\gamma=0.5$, and dotted for $\gamma=1$—without any markers, whereas the latter is plotted using red curves—solid for $\gamma=0$, dashed for $\gamma=0.5$, and dotted for $\gamma=1$—with circle $(\bullet)$ markers. The associated parameter setting can be found in the note of the figure. The left panel of the figure illustrates cases where bank 2’s robust optimal strategy guarantees its solvency regardless of $\gamma$. By contrast, the right panel depicts scenarios in which the proposed strategy leads to solvency when $\gamma=0$ and $0.5$, but default is inevitable when $\gamma=1$.
@@
In both panels, we observe that in the range of $x_{2}$ where the $\bullet$-marked red curves exceed $L_{2}=5$, the unmarked blue curves remain at $L_{2}=5$, which is consistent with Proposition 4.3(i). Additionally, whenever the $\bullet$-marked red curves are below $L_{2}=5$, they are positioned slightly beneath the unmarked blue curves, which confirms Proposition 4.3(ii), and the negligible gap between two curves demonstrates the reliability of our relaxed problem (9) at least in this example.
[…]
@@
A numerical comparison between the inner optimal values of (7) and (9) We use a two-bank system with Q(x)=10.2xQ(x)=1-0.2x, $L_{1}=L_{2}=5$, $z_{1}=1$, $z_{2}=0.5$, and $e_{2}=4.55$, under $\gamma=0$, $0.5$, and $1$. Also, we set $e_{1}=4.9$ and $e_{1}=4.85$ for the left and right panels, respectively.
@@
4.3 Analysis of the Robust Optimal Strategy with Permanent Price Impact
@@
Derivation of the robust optimal strategy.
When $\gamma=0$, (9) reduces to (2), leading to $x^{\gamma}_{i}=x^{*}_{i}=\bar{x}_{i}/2$. In contrast, when $\gamma>0$, early liquidation not only incurs an immediate price impact at $t=1$ but also depresses the asset price at $t=2$. Consequently, from bank $i$’s perspective, the worst-case scenario arises when other banks liquidate more aggressively at $t=1$. To counter this compounded impact at $t=1$, bank $i$ must employ a more prudent early liquidation strategy that depends on the permanency level $\gamma$, resulting in a downward deviation of the robust optimal strategy $x^{\gamma}_{i}$ from $x^{*}_{i}$ that solves (2). This is formalized in the following theorem. Furthermore, the theorem demonstrates that $x^{\gamma}_{i}$ is uniquely determined and admits a (semi-)closed-form expression for any $\gamma\in[0,1]$.
Theorem 4.4 (Robust Optimal Strategy with Permanent Price Impact)
@@
where $g_{i}^{\gamma}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q\left(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(x+s)\right)$. The equality in (12) holds due to the concave-convexity of $g_{i}^{\gamma}$ and the minimax theorem (Rockafellar2015, Theorem 36.3). Then, the first-order conditions for (12), $({\partial g_{i}^{\gamma}}/{\partial x})(x_{i},s_{i})=0$ and $({\partial g_{i}^{\gamma}}/{\partial s})(x_{i},s_{i})=0$, hold only at
@@
While the reasoning outlined above provides a straightforward path to constructing a solution to (9), the essence of Theorem 4.4 lies in establishing its uniqueness, thereby ruling out the possibility of any alternative solutions; see the proof of the theorem in Section 7 for more details. It is also worth highlighting that solving (11) is computationally tractable since $(\partial g_{i}^{\gamma}/\partial x)(\cdot,\bar{s}_{i})$ is strictly decreasing. Moreover, when the inverse demand function is linear—a common assumption in prior studies—the solution admits a closed form. Specifically, if $Q(x)=P-\alpha x$, then (11) can be written as $P-\alpha(2x+\bar{s}_{i})=P-\alpha((2-\gamma)\bar{x}_{i}+\gamma\bar{s}_{i}-2(1-\gamma)x)$, and thus,
@@
Structural dependence on the parameter $\gamma$.
The V-shaped dependence of the robust optimal strategy mentioned above is not limited to the specific choice of $Q(\cdot)$. Indeed, the following theorem demonstrates that, under a mild condition, this behavior holds for “all” inverse demand functions satisfying Assumption 4.1.
Theorem 4.5 (Impact of $\gamma$: A Typical Scenario)
@@
Firstly, we note that the condition $\gamma^{*}\geq 1/2$ is equivalent to $\bar{x}_{i}\leq\bar{s}_{i}\coloneqq\sum_{j\neq i}\bar{x}_{j}$, meaning that bank $i$’s contribution to the price impact is smaller than the aggregate contribution of all other banks. Thus, it is typically satisfied in practice. Secondly, as discussed earlier, the worst-case scenario for bank $i$ with $\gamma>\gamma^{*}$ corresponds to a situation in which all other banks collectively liquidate $\bar{s}_{i}$ units of illiquid assets exclusively at $t=1$. Therefore, Theorem 4.5 aligns with the intuition that, given $\bar{x}_{i}\leq\bar{s}_{i}$, a larger $\gamma$ leads to a sharper price decline at $t=2$. This prompts bank $i$ to increase sales at $t=1$ to avoid the anticipated future price distress. By combining this result with the analytical solution (10) for $x^{\gamma}_{i}$ when $\gamma\leq\gamma^{*}$, the aforementioned V-shaped dependence holds generally.
The left and right panels of Figure 4.3 depict a bank’s robust optimal strategy and its worst-case scenario on the aggregate early liquidation size of other banks, respectively, as functions of $\gamma$ when the inverse demand function $Q$ is nonlinear; specifically $Q(x)=\exp(-0.2x)$. In this example, $\gamma^{*}\approx 0.7$, satisfying the condition in Theorem 4.5. Consistent with our theoretical results, the left panel displays the aforementioned non-monotonic behavior of $x^{\gamma}_{i}$ with the tipping point $\gamma^{*}$ and the upper bound $\bar{x}_{i}/2$, and the right panel shows that under the worst-case scenario, higher $\gamma$ drives other banks to liquidate more at $t=1$ when $\gamma\leq\gamma^{*}$, and to liquidate exclusively in the first period once $\gamma>\gamma^{*}$.
[…]
A bank’s robust optimal strategy and its worst-case scenario on the aggregate early liquidation size of other banks under various levels of permanent price impact when $\gamma^{*}$ is large We use a four-bank system with $i=1$, $L_{1}=\cdots=L_{4}=6$, $z_{1}=\cdots=z_{4}=1$, $w_{1}=0.3$, $w_{2}=w_{3}=0.5$, and $w_{4}=0.8$. We set $Q(x)=\exp(-0.2x)$ and $s(x)=\mathop{\arg\min}_{s\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x,s)$, where $g_{i}^{\gamma}$ is defined as in (12).
@@
Proposition 4.6 (Impact of $x^{\gamma}_{i}$: An Extreme Scenario)
@@
In Figure 4.3, we conduct an experiment similar to that in Figure 4.3, but with $\gamma^{*}\approx 0.11$. In this case, the non-monotonicity vanishes; instead, $x^{\gamma}_{i}$ decreases monotonically with $\gamma$, validating Theorem 4.6. The right panel of the figure mirrors the pattern in that of Figure 4.3: increasing with $\gamma$ until $\gamma\leq\gamma^{*}$ and plateauing at $\bar{s}_{i}$ thereafter.
[…]

### 17 Mar 2026 v2

@@ 4.1.  Model Setup
Balance sheet information. We consider a financial system comprised of $m$ banks, indexed by $1,\dots,m$, that operates over two periods ($t=1,2$). Each bank $i$’s balance sheet consists of the following four components:
@@
  • $z_{i}\geq 0$ represents the “volume” of its illiquid assets, with a nominal unit price of PP; [^3] We adopt the standard assumption in the literature that all banks hold the same type of illiquid assets, which serves as an approximation of multiple correlated assets (Weber and Weske, 2017; Braouezec and Wagalath, 2019).
@@
We assume that these components remain constant throughout the two periods and that liquid assets (e.g., cash, central bank reserves, and short-term government securities) can be sold and transferred at their nominal value without any cost. To exclude trivial cases where banks are not required to liquidate their illiquid assets, it is assumed that $e_{i}<L_{i}$ for all $i\in[m]$. The associated balance sheet is described in Table 1.
Table 1: Initial balance sheet of bank $i$ before liquidation
Assets Liabilities and Net Worth
Liquid assets $e_{i}$ Liabilities $L_{i}$
Illiquid assets $z_{i}$ with price $P$ Net worth $w_{i}$
@@
Illiquid asset price modeling. Illiquid assets (e.g., loans, mortgages, and real estate holdings) are, however, difficult to sell without a significant loss of value. In the related literature (Cifuentes et al., 2005; Chen et al., 2016), this loss is typically modeled by the so-called inverse demand function $Q:[0,\sum_{j=1}^{m}z_{j}]\rightarrow\mathbb{R}_{+}$, a decreasing function that maps the aggregate supply of illiquid assets across all banks to their price. Throughout this paper, we impose several regularity conditions on the function $Q$, described below:
Assumption 1.
The inverse demand function $Q$ is twice continuously differentiable and satisfies the following five conditions: (i) $Q(0)=P$; (ii) $Q^{\prime}(x)<0$; (iii) $(xQ(x))^{\prime}>0$; (iv) $Q^{\prime\prime}(x)\geq 0$; and (v) $(xQ^{\prime}(x))^{\prime}<0$.
The first condition states that illiquid assets are valued at their nominal price in the absence of sales. Conditions (ii) and (iii) reflect the observation that as the supply of illiquid assets increases, their price decreases while the total value of these assets in the market grows. The fourth condition imposes a diminishing marginal effect on the price of illiquid assets with respect to their size. Condition (v) ensures that the total value of illiquid assets in the market is concave in their size. Note that the above conditions (or their variants) are commonly used in the literature (Amini et al., 2016; Bichuch and Feinstein, 2019). Examples of inverse demand functions satisfying Assumption 1 include Q(x)=PaxQ(x)=P-ax with $0<a<P/(2\sum_{k=1}^{m}z_{k})$, $Q(x)=P\exp(-ax)$ with $0<a<1/\sum_{k=1}^{m}z_{k}$, and $Q(x)=P(\beta x+1)^{-a}$ with $0<a<1/(\beta\sum_{k=1}^{m}z_{k})$ and $\beta>0$.
@@
Early liquidation and permanent price impact. We assume that all banks clear their debts at the end of the time horizon (i.e., $t=2$), consistent with prior studies (e.g., Cifuentes et al., 2005; Chen et al., 2016; Amini et al., 2016; Ghamami et al., 2022). As alluded to earlier, a salient feature of our modeling framework lies in the option for strategic early liquidation at $t=1$ to mitigate liquidation losses and to sell a smaller volume of illiquid assets, which is in stark contrast to most existing models that allow liquidation only at clearing. Suppose that each bank $i$ chooses to sell off $x_{i}$ units of illiquid assets at $t=1$. Then, $Q(\sum_{k=1}^{m}x_{k})$ becomes the market price of illiquid assets at $t=1$ and is strictly smaller than the nominal price $P$ by Assumption 1.
@@
To account for permanent price impact, we consider a situation where the price impact generated at $t=1$ persists into $t=2$. Specifically, the asset price at clearing is governed by another inverse demand function of the asset supply at $t=2$, defined as $u\mapsto Q(\gamma\sum_{k=1}^{m}x_{k}+u)$, where the parameter $\gamma\in[0,1]$ represents the proportion of the early liquidation volume that continues to depress prices at $t=2$. That is, a larger $\gamma$ implies slower price recovery. The limiting case $\gamma=0$ corresponds to fully transient impact, where the initial price at $t=2$ recovers to the nominal price $P$, as in the example in Section 3. Conversely, $\gamma=1$ signifies fully permanent price impact, where the opening price at $t=2$ remains at the post-liquidation price at $t=1$; this latter case coincides with the model in Ghamami et al. (2022). While existing research commonly focuses on either of these two extreme cases, our approach generalizes the framework of Ghamami et al. (2022) by allowing for “partially” permanent price impact via the parameter $\gamma$.
@@
Clearing process at $t=2$. Following the early liquidation at $t=1$, each bank $i\in[m]$ retains $z_{i}-x_{i}$ units of illiquid assets, while its liquid asset value becomes $e_{i}+x_{i}Q(\sum_{k=1}^{m}x_{k})$. At the clearing stage, given a price $p$, each bank $i$ attempts to cover its liabilities by selling $p^{-1}(L_{i}-e_{i}-x_{i}Q(\sum_{k=1}^{m}x_{k}))^{+}$ units of illiquid assets. If the bank cannot meet the obligations, it defaults and liquidates all remaining $z_{i}-x_{i}$ units of assets. Thus, the quantity sold by bank $i$ is $\{p^{-1}(L_{i}-e_{i}-x_{i}Q(\sum_{k=1}^{m}x_{k}))^{+}\}\wedge(z_{i}-x_{i})$ for all $i\in[m]$. Consequently, the market price $p^{\gamma}_{2}$ of illiquid assets at $t=2$ is determined by the following fixed-point equation
@@
Note that a straightforward extension of Lemma 3 in Amini et al. (2016) ensures the uniqueness of $p^{\gamma}_{2}$. We denote each bank $i$’s liquidation size at $t=2$ by
@@
Remark 1.
Permanent price impact could also be modeled using more sophisticated dynamics, such as, the Volume Weighted Average Price (VWAP) approach investigated in  Banerjee and Feinstein (2021). Unlike static snapshots, this method captures the entire liquidation history by defining the average asset price through an integral of the inverse demand function. However, the technical analysis based on the VWAP framework introduces significant mathematical complexity due to the integration terms; therefore, we retain our current setup to preserve analytical tractability. We anticipate that the qualitative insights derived from our findings would persist under the VWAP framework.
4.2.  Main Problem
@@
Robustification against other banks’ actions. As shown in (6), each bank’s liquid asset value at clearing depends on the volume of illiquid assets sold by other banks at $t=1$, which is often unknown. Hence, the solution to the following problem yields bank $i$’s optimal early liquidation decision under the worst-case strategies of other banks:
@@
Relaxation of liquidation size at clearing. Invoking Lemma 1, we circumvent the aforementioned intractability of (7) by relaxing the constraint on each bank $j$’s liquidation size at clearing. Specifically, we allow it to take any value in the interval $0,\bar{x}_{j}-x_{j}]$ rather than restricting it to $\psi_{j}({\bf x})$ for each $j\in[m]$. Under this relaxation, the maximin problem in (7) can be converted into the following form:
@@
For any feasible solution $({\bf x},{\bf y})$ of (8), Assumption 1 indicates that
@@
Validity of the formulation. Before presenting the analysis of our problem (9), let us discuss its practical validity to address potential concerns about the implication of optimizing $\Phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$. Specifically, the following result provides insights into how the relaxation affects the liquidation strategy, or equivalently, the relationship between the solutions to (7) and (9)
Proposition 1 (Validity of (9)).
@@
By the definitions of $\bar{x}_{i}$ and $\psi^{\gamma}_{j}({\bf x})$, it is easy to check that $\phi^{\gamma}_{i}(x_{i},\bar{{\bf x}}_{-i})\leq L_{i}$ for any $x_{i}\in[0,\bar{x}_{i}]$; see the proof of Proposition 1 for more details. [^5] From a practical perspective, this inequality holds since each bank’s liquid asset value at clearing is characterized as either exactly meeting the payment obligation or failing to do so. Thus, the first statement of Proposition 1 shows that if the optimal value of (9) is greater than or equal to $L_{i}$, then bank $i$’s robust optimal strategy, $x^{\gamma}_{i}$, is also an optimal solution to the outer problem of (7). Furthermore, the second statement of the theorem suggests that, even if the optimal value of (9) is smaller than $L_{i}$, bank $i$’s robust optimal strategy still maximizes a lower bound for the worst-case value of its liquid assets at clearing. Accordingly, both statements indicate that the robust optimal strategy obtained by solving (9) can serve as a conservative approximation and a reliable surrogate for the true optimal decision that solves (7). This answers our second research question (Q2), introduced in Section 3, for the case of permanent price impact.
@@
In Figure 1, we numerically validate the results in Proposition 1 by comparing the inner optimal values $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$ and $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\Phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$ across different values of $x_{i}$ in a two-bank system with $m=i=2$ for varying levels of permanent price impact: $\gamma=$ 0, $0.5$, and $1$. The former optimal value is plotted using blue curves—solid for $\gamma=0$, dashed for $\gamma=0.5$, and dotted for $\gamma=1$—without any markers, whereas the latter is plotted using red curves—solid for $\gamma=0$, dashed for $\gamma=0.5$, and dotted for $\gamma=1$—with circle $(\bullet)$ markers. The associated parameter setting can be found in the note of the figure. The left panel of the figure illustrates cases where bank 2’s robust optimal strategy guarantees its solvency regardless of $\gamma$. By contrast, the right panel depicts scenarios in which the proposed strategy leads to solvency when $\gamma=0$ and $0.5$, but default is inevitable when $\gamma=1$.
@@
In both panels, we observe that in the range of $x_{2}$ where the $\bullet$-marked red curves exceed $L_{2}=5$, the unmarked blue curves remain at $L_{2}=5$, which is consistent with Proposition 1(i). Additionally, whenever the $\bullet$-marked red curves are below $L_{2}=5$, they are positioned slightly beneath the unmarked blue curves, which confirms Proposition 1(ii), and the negligible gap between two curves demonstrates the reliability of our relaxed problem (9) at least in this example.
@@
Figure 1: A numerical comparison between the inner optimal values of (7) and (9). We use a two-bank system with Q(x)=10.2xQ(x)=1-0.2x, $L_{1}=L_{2}=5$, $z_{1}=1$, $z_{2}=0.5$, and $e_{2}=4.55$, under $\gamma=0$, $0.5$, and $1$. Also, we set $e_{1}=4.9$ and $e_{1}=4.85$ for the left and right panels, respectively.
@@
4.3.  Analysis of the Robust Optimal Strategy with Permanent Price Impact
@@
Derivation of the robust optimal strategy. When $\gamma=0$, (9) reduces to (2), leading to $x^{\gamma}_{i}=x^{*}_{i}=\bar{x}_{i}/2$. In contrast, when $\gamma>0$, early liquidation not only incurs an immediate price impact at $t=1$ but also depresses the asset price at $t=2$. Consequently, from bank $i$’s perspective, the worst-case scenario arises when other banks liquidate more aggressively at $t=1$. To counter this compounded impact at $t=1$, bank $i$ must employ a more prudent early liquidation strategy that depends on the permanency level $\gamma$, resulting in a downward deviation of the robust optimal strategy $x^{\gamma}_{i}$ from $x^{*}_{i}$ that solves (2). This is formalized in the following theorem. Furthermore, the theorem demonstrates that $x^{\gamma}_{i}$ is uniquely determined and admits a (semi-)closed-form expression for any $\gamma\in[0,1]$.
Theorem 1 (Robust Optimal Strategy with Permanent Price Impact).
@@
where $g_{i}^{\gamma}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q\left(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(x+s)\right)$. The equality in (12) holds due to the concave-convexity of $g_{i}^{\gamma}$ and the minimax theorem (Rockafellar, 2015, Theorem 36.3). Then, the first-order conditions for (12), $({\partial g_{i}^{\gamma}}/{\partial x})(x_{i},s_{i})=0$ and $({\partial g_{i}^{\gamma}}/{\partial s})(x_{i},s_{i})=0$, hold only at
@@
While the reasoning outlined above provides a straightforward path to constructing a solution to (9), the essence of Theorem 1 lies in establishing its uniqueness, thereby ruling out the possibility of any alternative solutions; see the proof of the theorem in Section A for more details. It is also worth highlighting that solving (11) is computationally tractable since $(\partial g_{i}^{\gamma}/\partial x)(\cdot,\bar{s}_{i})$ is strictly decreasing. Moreover, when the inverse demand function is linear—a common assumption in prior studies—the solution admits a closed form. Specifically, if $Q(x)=P-\alpha x$, then (11) can be written as $P-\alpha(2x+\bar{s}_{i})=P-\alpha((2-\gamma)\bar{x}_{i}+\gamma\bar{s}_{i}-2(1-\gamma)x)$, and thus,
@@
Structural dependence on the parameter $\gamma$. The V-shaped dependence of the robust optimal strategy mentioned above is not limited to the specific choice of $Q(\cdot)$. Indeed, the following theorem demonstrates that, under a mild condition, this behavior holds for “all” inverse demand functions satisfying Assumption 1.
Theorem 2 (Impact of $\gamma$: A Typical Scenario).
@@
Firstly, we note that the condition $\gamma^{*}\geq 1/2$ is equivalent to $\bar{x}_{i}\leq\bar{s}_{i}\coloneqq\sum_{j\neq i}\bar{x}_{j}$, meaning that bank $i$’s contribution to the price impact is smaller than the aggregate contribution of all other banks. Thus, it is typically satisfied in practice. Secondly, as discussed earlier, the worst-case scenario for bank $i$ with $\gamma>\gamma^{*}$ corresponds to a situation in which all other banks collectively liquidate $\bar{s}_{i}$ units of illiquid assets exclusively at $t=1$. Therefore, Theorem 2 aligns with the intuition that, given $\bar{x}_{i}\leq\bar{s}_{i}$, a larger $\gamma$ leads to a sharper price decline at $t=2$. This prompts bank $i$ to increase sales at $t=1$ to avoid the anticipated future price distress. By combining this result with the analytical solution (10) for $x^{\gamma}_{i}$ when $\gamma\leq\gamma^{*}$, the aforementioned V-shaped dependence holds generally.
The left and right panels of Figure 2 depict a bank’s robust optimal strategy and its worst-case scenario on the aggregate early liquidation size of other banks, respectively, as functions of $\gamma$ when the inverse demand function $Q$ is nonlinear; specifically $Q(x)=\exp(-0.2x)$. In this example, $\gamma^{*}\approx 0.7$, satisfying the condition in Theorem 2. Consistent with our theoretical results, the left panel displays the aforementioned non-monotonic behavior of $x^{\gamma}_{i}$ with the tipping point $\gamma^{*}$ and the upper bound $\bar{x}_{i}/2$, and the right panel shows that under the worst-case scenario, higher $\gamma$ drives other banks to liquidate more at $t=1$ when $\gamma\leq\gamma^{*}$, and to liquidate exclusively in the first period once $\gamma>\gamma^{*}$.
Figure 2: A bank’s robust optimal strategy and its worst-case scenario on the aggregate early liquidation size of other banks under various levels of permanent price impact when $\gamma^{*}$ is large. We use a four-bank system with $i=1$, $L_{1}=\cdots=L_{4}=6$, $z_{1}=\cdots=z_{4}=1$, $w_{1}=0.3$, $w_{2}=w_{3}=0.5$, and $w_{4}=0.8$. We set $Q(x)=\exp(-0.2x)$ and $s(x)=\mathop{\arg\min}_{s\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x,s)$, where $g_{i}^{\gamma}$ is defined as in (12).
@@
Proposition 2 (Impact of $x^{\gamma}_{i}$: An Extreme Scenario).
@@
In Figure 3, we conduct an experiment similar to that in Figure 2, but with $\gamma^{*}\approx 0.11$. In this case, the non-monotonicity vanishes; instead, $x^{\gamma}_{i}$ decreases monotonically with $\gamma$, validating Theorem 2. The right panel of the figure mirrors the pattern in that of Figure 2: increasing with $\gamma$ until $\gamma\leq\gamma^{*}$ and plateauing at $\bar{s}_{i}$ thereafter.